dalelewis:

Find the equation of the axis of symmetry and the coordinates of the vertex of the graph of the function y = 4x2 + 5x – 1

jhonyy9:

any idea ?

charge69:

Formula for AOS is x=-b/2a B=5 A=4 -5/2(4) -5/8 X=-5/8 is the aos